Subject: RAITH 150 Start up status update 1730 hrs. WED. 1-7-2003
From: James Conway <jwc@snf.stanford.edu>
Date: Wed, 07 Jan 2004 17:38:36 -0800
Wed, 07 Jan 2004 17:38:36 -0800
Hello Folks,

Startup Status:
We are still working to bring the system back up after shut down.
We have encountered several minor problems, which have now been 
resolved, but my original time estimate has slipped back by at least 
another day.   We are hoping now to allow access to the system Friday 
with some opportunities to help us test Thursday afternoon and 
overnight. Full update tomorrow at Noon.

RAITH:
Today I have completed the X-Y stage calibrations and the stage is 
working superbly.
The stage leveling control section is next...

LEO and FE-Gun:
This afternoon we had 'first light' on the FE-Gun and it will require 
overnight to stabilize.
Thursday AM we plan to realign the column and adjust and calibrate the 
three magnification amplifiers and perform the centering calibrations 
for the aperture align and stigmation amplifiers to the deflection plate 
section.

Once we have completed these task if the beam current and imaging 
resolution is acceptable we will start with qualification of the system 
for line width, resolution, and stitching. Note that SEM measurements 
made at the last day of the year DID NOT display suitable imaging 
resolution and we are evaluating the FE-tip condition with respect to 
the beam diameter.
(read: Beam spot size)

Both system's data and image files were back up today and new emergency 
recovery disk were generated.

What does this mean for You?
Please don't expect access to the system until we are up and have 
established new baselines for the system.  I am hoping we can have this 
finished by late Friday.  There possibly may be access for people 
willing to help test the system on Thursday afternoon and overnight for 
RAITH Champions and experienced users whom are reserved on the CORAL 
schedule.
Please see me directly if you have questions.

Thank you for your support!

James Conway


Hello Folks,

Startup Status:
We are still working to bring the system back up after shut down.
We have encountered several minor problems, which have now been resolved, but my original time estimate has slipped back by at least another day.   We are hoping now to allow access to the system Friday with some opportunities to help us test Thursday afternoon and overnight. Full update tomorrow at Noon.

RAITH:

Today I have completed the X-Y stage calibrations and the stage is working superbly.
The stage leveling control section is next...

LEO and FE-Gun:
This afternoon we had 'first light' on the FE-Gun and it will require overnight to stabilize.
Thursday AM we plan to realign the column and adjust and calibrate the three magnification amplifiers and perform the centering calibrations for the aperture align and stigmation amplifiers to the deflection plate section.

Once we have completed these task if the beam current and imaging resolution is acceptable we will start with qualification of the system for line width, resolution, and stitching. Note that SEM measurements made at the last day of the year DID NOT display suitable imaging resolution and we are evaluating the FE-tip condition with respect to the beam diameter.
(read: Beam spot size)

Both system's data and image files were back up today and new emergency recovery disk were generated.

What does this mean for You?

Please don't expect access to the system until we are up and have established new baselines for the system.  I am hoping we can have this finished by late Friday.  There possibly may be access for people willing to help test the system on Thursday afternoon and overnight for RAITH Champions and experienced users whom are reserved on the CORAL schedule.
Please see me directly if you have questions.

Thank you for your support!

James Conway